June 20, 2012 Tracking CDC/FDC

From GlueXWiki
Jump to: navigation, search

Meeting Time and Place

Wednesday June 20, 2012 at 11:00am At Jefferson Lab, the meeting will be held in F326


To connect from the outside, please use ESNET

  1. ) ESNET: 8542553
  2. )To connect by telephone, dial:
  • You can look up toll-free number at http://www.readytalk.com/intl
  • US and Canada: (866)740-1260 (toll free)
  • International: (303)248-0285 (toll call)
  • enter access code followed by the # sign: 3421244#

Hardware Agenda

  1. Announcements
  2. FDC Update
  3. CDC Update
    1. Construction Update
    2. Cosmic Calibration of the Chamber
  4. TDR Update

Tracking Software Agenda

  1. Announcements

Hardware Minutes

Participants: Naomi, Beni, Simon, Lubomir, Tim, Mark

  1. Announcements: Curtis is in Italy.
  2. FDC Update
    1. Continue with the stacking of the fourth package. The last three cells will be installed by the end of the week.
    2. Testing first production wire frames, that will be spares. First one can't be used, have to build a new one. The second one was fixed (a spot from a chemical was removed) but still showing high current. For these frames we used Epolite (epoxy) that is highly susceptible to humidity, so we hope after flushing with gas the frame will work.
    3. Working on pre-amp assignment to the electronics. The cards with highest rates at the center of the chambers are combined with those at the periphery to equalize the loads on the modules.
    4. Data taken with a 55Fe source and fADC125 using a strip for triggering show no long tails after the main pulse; this must be a feature only of the CDC detector/pre-amps.
  3. CDC Update
    1. Construction:
      Cdc insects1.jpg
      Cdc insects2.jpg
      The team are mid-way through row D, see photos (the yellow labels are temporary). We extended the choice of wire lengths to include 4.25" (yellow heatshrink) and 4.75" (white). We have paper documentation containing the map for each board, with any changes in pin to pad assignments noted. The technicians are also maintaining an electronic file with the pin to pad assignments; the assignments are sometimes altered deliberately, for example, if 2 wires to neighboring pins on the same board cross each other, then uncrossing them makes it easier to fit another wire (for a board in the next ring out) between them later on. There are a number of checks made when each transition board has been completed.
      1. A visual check is made by the technician who did not solder the board, looking at pin to pad matching and soldering and that the conductive rubber completely covers the crimp pin. A gentle wiggle-test is also done to make sure that the soldering to the board is sound. These checks are repeated later by Gary or Naomi.
      2. The completed boards are tested by applying 1500V to a temporarily inserted HVB. Most boards have shown near-zero current. A few boards have drawn small current (few tens of nA) and this has reduced to near-zero after leaving the voltage on the board for a while (30 mins to a few hours).
    2. Cabling: Beni said that Fernando was happy with the cable plan supplied so far, he needs to think more about how to assign the HVBs to the fADCs, maybe interleave each higher-rate board from the inner rows between 2 lower-rate boards from the outermost.

Tracking Software Minutes

Participants: Naomi, Paul, Beni, Simon, Lubomir, Tim, Mark, Dave

  1. Meeting Adjourned.